# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ Appendix C Full Results on D-all and A-all Datasets
Tables 12, 13 and 14 show results for all models on D-all and A-all datasets. Among models with SFT Gemma2 2B IT is the best on D-all, while on A-all the best are encoder-based models – RuRoBERTa and BERT. While in general prompting performs worse than SFT on D-all, the best model (Gemma 2 9B IT) shows comparable performance with AutoML models. On the contrary, on A-all prompting shows better results than SFT. The best models in various prompting settings are Qwen2 7B IT, SaigaLlama3 8B, and Gemma2 2B IT.
@@
| D-all | Gemma2 2B IT | 97.141.04 | 92.330.94 | 94.670.72 | 52.986.33 | 76.446.36 | 62.335.6 | 78.53.13 | ||||||||
| Gemma2 9B IT | 97.650.82 | 91.581.24 | 94.510.68 | 47.628.42 | 77.866.23 | 58.696.91 | 76.63.77 | |||||||||
| Qwen2 7B IT | 97.030.96 | 91.090.64 | 93.960.17 | 44.644.94 | 72.975.26 | 54.982.77 | 74.471.36 | |||||||||
| SaigaLlama3 8B | 96.731.49 | 90.881.0 | 93.711.01 | 43.456.65 | 70.1610.04 | 53.417.35 | 73.564.13 | |||||||||
| Vikhr 7B IT 0.4 | 96.831.2 | 90.461.23 | 93.531.01 | 40.488.16 | 68.8611.28 | 50.738.61 | 72.134.79 | |||||||||
| Vikhr 7B IT 5.4 | 92.021.28 | 90.181.16 | 91.091.16 | 41.677.04 | 47.327.9 | 44.287.35 | 67.694.24 | |||||||||
| VikhrGemma 2B IT | 97.441.68 | 91.830.8 | 94.540.66 | 49.45.98 | 79.2511.48 | 60.043.8 | 77.292.09 | |||||||||
| RuRoBERTa | 92.480.63 | 95.51.16 | 93.960.42 | 68.195.05 | 54.764.45 | 60.482.31 | 77.221.27 | |||||||||
| RuBioRoBERTa | 90.742.52 | 96.931.66 | 93.691.04 | 58.526.99 | 41.6719.2 | 48.6122.33 | 71.1511.6 | |||||||||
| RuBERT | 92.260.52 | 96.221.56 | 94.190.68 | 72.1210.1 | 52.983.81 | 60.62.87 | 77.391.72 | |||||||||
| BERT | 90.163.31 | 96.631.69 | 93.221.43 | 50.5228.15 | 37.524.38 | 42.526.53 | 67.8613.9 | |||||||||
| Linguistic features | 88.10.6 | 88.13.4 | 88.01.6 | 51.65.8 | 50.74.1 | 50.82.6 | 69.41.9 | |||||||||
| TF-IDF | 89.51.0 | 84.21.0 | 86.80.6 | 47.51.8 | 59.14.2 | 52.62.5 | 69.71.5 | |||||||||
| A-all | Gemma2 2B IT | 57.094.28 | 57.613.21 | 57.273.15 | 51.116.0 | 50.613.62 | 50.774.48 | 54.023.39 | ||||||||
| Gemma2 9B IT | 56.324.04 | 55.441.46 | 55.832.54 | 47.563.14 | 48.471.77 | 47.951.96 | 51.891.63 | |||||||||
| Qwen2 7B IT | 56.133.95 | 56.431.57 | 56.232.59 | 49.782.85 | 49.512.04 | 49.591.86 | 52.911.79 | |||||||||
| SaigaLlama3 8B | 56.324.83 | 58.561.32 | 57.332.79 | 53.783.74 | 51.592.03 | 52.561.83 | 54.951.57 | |||||||||
| Vikhr 7B IT 0.4 | 56.132.43 | 57.083.28 | 56.582.68 | 50.894.95 | 49.923.4 | 50.384.09 | 53.483.32 | |||||||||
| Vikhr 7B IT 5.4 | 55.367.22 | 56.892.4 | 55.954.71 | 51.564.79 | 50.13.07 | 50.652.68 | 53.32.82 | |||||||||
| VikhrGemma 2B IT | 53.262.94 | 54.792.65 | 53.972.34 | 48.894.91 | 47.342.63 | 48.063.59 | 51.012.62 | |||||||||
| RuRoBERTa | 61.533.69 | 60.1513.8 | 59.684.47 | 54.390.97 | 54.4418.08 | 52.2711.87 | 55.973.88 | |||||||||
@@
| RuBERT | 57.181.93 | 51.535.61 | 54.093.67 | 49.691.98 | 55.334.34 | 52.252.28 | 53.172.1 | |||||||||
| BERT | 59.441.74 | 46.933.66 | 52.332.1 | 50.430.98 | 62.674.81 | 55.822.21 | 54.081.08 | |||||||||
| Linguistic features | 53.42.9 | 39.719.3 | 42.017.9 | 46.12.4 | 60.220.4 | 51.07.8 | 46.56.2 | |||||||||
| TF-IDF | 51.80.7 | 40.015.6 | 43.310.8 | 44.32.1 | 56.417.8 | 48.68.5 | 46.02.4 |
@@
| D-all | Gemma2 2B IT | 5-shot MMLU | 7.36 | 100.0 | 13.71 | 100.0 | 15.64 | 27.05 | 20.38 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 30.06 | 100.0 | 46.23 | 100.0 | 19.72 | 32.94 | 39.58 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 97.55 | 89.33 | 93.26 | 32.14 | 69.23 | 43.90 | 68.58 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 68.1 | 94.87 | 79.29 | 78.57 | 29.73 | 43.14 | 61.21 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 22.09 | 94.74 | 35.82 | 92.86 | 16.99 | 28.73 | 32.28 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 52.15 | 95.51 | 67.46 | 85.71 | 23.53 | 36.92 | 52.19 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.61 | 100.0 | 1.22 | 100.0 | 14.74 | 25.69 | 13.45 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 41.1 | 93.06 | 57.02 | 82.14 | 19.33 | 31.29 | 44.16 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 89.57 | 88.48 | 89.02 | 32.14 | 34.62 | 33.33 | 61.18 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 2.45 | 100.0 | 4.79 | 100.0 | 14.97 | 26.05 | 15.42 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 57.06 | 96.88 | 71.81 | 89.29 | 26.32 | 40.65 | 56.23 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 31.29 | 91.07 | 46.58 | 82.14 | 17.04 | 28.22 | 37.4 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 38.04 | 84.93 | 52.54 | 60.71 | 14.41 | 23.29 | 37.92 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 98.77 | 85.19 | 91.48 | 0.0 | 0.0 | 0.0 | 45.74 | |||||||||
| A-all | Gemma2 2B IT | 5-shot MMLU | 10.34 | 64.29 | 17.82 | 93.33 | 47.3 | 62.78 | 40.3 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 11.49 | 66.67 | 19.61 | 93.33 | 47.62 | 63.06 | 41.34 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 73.56 | 57.14 | 64.32 | 36.0 | 54.0 | 43.2 | 53.76 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 39.08 | 65.38 | 48.92 | 76.0 | 51.82 | 61.62 | 55.27 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 56.32 | 62.82 | 59.39 | 61.33 | 54.76 | 57.86 | 58.63 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 62.07 | 62.79 | 62.43 | 57.33 | 56.58 | 56.95 | 59.69 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 10.34 | 69.23 | 18.0 | 94.67 | 47.65 | 63.39 | 40.7 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 34.48 | 69.77 | 46.15 | 82.67 | 52.1 | 63.92 | 55.04 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 96.55 | 53.85 | 69.14 | 4.0 | 50.0 | 7.41 | 38.27 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 22.99 | 62.5 | 33.61 | 84.0 | 48.46 | 61.46 | 47.54 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 13.79 | 70.59 | 23.08 | 93.33 | 48.28 | 63.64 | 43.36 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 33.33 | 59.18 | 42.65 | 73.33 | 48.67 | 58.51 | 50.58 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 94.25 | 54.67 | 69.2 | 9.33 | 58.33 | 16.09 | 42.65 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 98.85 | 53.42 | 69.35 | 0.0 | 0.0 | 0.0 | 34.68 |

### v3

@@ Appendix C Full Results on D-all and A-all Datasets
Tables˜14, 15 and 16 show results for all models on D-all and A-all datasets. Among models with SFT Gemma2 2B IT is the best on D-all, while on A-all the best are encoder-based models – RuRoBERTa and BERT. While in general prompting performs worse than SFT on D-all, the best model (Gemma 2 9B IT) shows comparable performance with AutoML models. On the contrary, prompting on the A-all shows better results than SFT. The best models in various prompting settings are Qwen2 7B IT, SaigaLlama3 8B, and Gemma2 2B IT.
@@
| D-all | Gemma2 2B IT | 92.33 0.94 | 97.14 1.04 | 94.67 0.72 | 76.44 6.36 | 52.98 6.33 | 62.33 5.60 | 78.50 3.13 | ||||||||
| Gemma2 9B IT | 91.58 1.24 | 97.65 0.82 | 94.51 0.68 | 77.86 6.23 | 47.62 8.42 | 58.69 6.91 | 76.60 3.77 | |||||||||
| Qwen2 7B IT | 91.09 0.64 | 97.03 0.96 | 93.96 0.17 | 72.97 5.26 | 44.64 4.94 | 54.98 2.77 | 74.47 1.36 | |||||||||
| SaigaLlama3 8B | 90.88 1.00 | 96.73 1.49 | 93.71 1.01 | 70.16 10.04 | 43.45 6.65 | 53.41 7.35 | 73.56 4.13 | |||||||||
| Vikhr 7B IT 0.4 | 90.46 1.23 | 96.83 1.20 | 93.53 1.01 | 68.86 11.28 | 40.48 8.16 | 50.73 8.61 | 72.13 4.79 | |||||||||
| Vikhr 7B IT 5.4 | 90.18 1.16 | 92.02 1.28 | 91.09 1.16 | 47.32 7.90 | 41.67 7.04 | 44.28 7.35 | 67.69 4.24 | |||||||||
| VikhrGemma 2B IT | 91.83 0.80 | 97.44 1.68 | 94.54 0.66 | 79.25 11.48 | 49.40 5.98 | 60.04 3.80 | 77.29 2.09 | |||||||||
| RuRoBERTa | 92.48 0.63 | 95.50 1.16 | 93.96 0.42 | 68.19 5.05 | 54.76 4.45 | 60.48 2.31 | 77.22 1.27 | |||||||||
| RuBioRoBERTa | 90.74 2.52 | 96.93 1.66 | 93.69 1.04 | 58.50 26.99 | 41.67 19.20 | 48.61 22.33 | 71.15 11.60 | |||||||||
| RuBERT | 92.26 0.52 | 96.22 1.56 | 94.19 0.68 | 72.12 10.10 | 52.98 3.81 | 60.60 2.87 | 77.39 1.72 | |||||||||
| BERT | 90.16 3.31 | 96.63 1.69 | 93.22 1.43 | 50.52 28.15 | 37.50 24.38 | 42.50 26.53 | 67.86 13.90 | |||||||||
| Linguistic features | 88.10 0.60 | 88.10 3.40 | 88.00 1.60 | 51.60 5.80 | 50.70 4.10 | 50.80 2.60 | 69.40 1.90 | |||||||||
| TF-IDF | 89.50 1.00 | 84.20 1.00 | 86.80 0.60 | 47.50 1.80 | 59.10 4.20 | 52.60 2.50 | 69.70 1.50 | |||||||||
| A-all | Gemma2 2B IT | 57.61 3.21 | 57.09 4.28 | 57.27 3.15 | 50.61 3.62 | 51.11 6.00 | 50.77 4.48 | 54.02 3.39 | ||||||||
| Gemma2 9B IT | 55.44 1.46 | 56.32 4.04 | 55.83 2.54 | 48.47 1.77 | 47.56 3.14 | 47.95 1.96 | 51.89 1.63 | |||||||||
| Qwen2 7B IT | 56.43 1.57 | 56.13 3.95 | 56.23 2.59 | 49.51 2.04 | 49.78 2.85 | 49.59 1.86 | 52.91 1.79 | |||||||||
| SaigaLlama3 8B | 58.56 1.32 | 56.32 4.83 | 57.33 2.79 | 51.59 2.03 | 53.78 3.74 | 52.56 1.83 | 54.95 1.57 | |||||||||
| Vikhr 7B IT 0.4 | 57.08 3.28 | 56.13 2.43 | 56.58 2.68 | 49.92 3.40 | 50.89 4.95 | 50.38 4.09 | 53.48 3.32 | |||||||||
| Vikhr 7B IT 5.4 | 56.89 2.40 | 55.36 7.22 | 55.95 4.71 | 50.10 3.07 | 51.56 4.79 | 50.65 2.68 | 53.30 2.82 | |||||||||
| VikhrGemma 2B IT | 54.79 2.65 | 53.26 2.94 | 53.97 2.34 | 47.34 2.63 | 48.89 4.91 | 48.06 3.59 | 51.01 2.62 | |||||||||
| RuRoBERTa | 61.53 3.69 | 60.15 13.80 | 59.68 4.47 | 54.39 0.97 | 54.44 18.08 | 52.27 11.87 | 55.97 3.88 | |||||||||
@@
| RuBERT | 57.18 1.93 | 51.53 5.61 | 54.09 3.67 | 49.69 1.98 | 55.33 4.34 | 52.25 2.28 | 53.17 2.10 | |||||||||
| BERT | 59.44 1.74 | 46.93 3.66 | 52.33 2.10 | 50.43 0.98 | 62.67 4.81 | 55.82 2.21 | 54.08 1.08 | |||||||||
| Linguistic features | 53.40 2.90 | 39.70 19.30 | 42.00 17.90 | 46.10 2.40 | 60.20 20.40 | 51.00 7.80 | 46.50 6.20 | |||||||||
| TF-IDF | 51.80 0.70 | 40.00 15.60 | 43.30 10.80 | 44.30 2.10 | 56.40 17.80 | 48.60 8.50 | 46.00 2.40 |
@@
| D-all | Gemma2 2B IT | 5-shot MMLU | 100.00 0.00 | 7.36 0.00 | 13.71 0.00 | 15.64 0.00 | 100.00 0.00 | 27.05 0.00 | 20.38 0.00 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 100.00 0.00 | 30.55 0.25 | 46.80 0.29 | 19.83 0.06 | 100.00 0.00 | 33.10 0.08 | 39.95 0.18 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 88.83 0.25 | 96.56 0.49 | 92.53 0.36 | 59.56 4.84 | 29.29 1.43 | 39.26 2.32 | 65.90 1.34 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 94.87 0.00 | 68.10 0.00 | 79.29 0.00 | 29.73 0.00 | 78.57 0.00 | 43.14 0.00 | 61.21 0.00 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 94.74 0.00 | 22.09 0.00 | 35.82 0.00 | 16.99 0.00 | 92.86 0.00 | 28.73 0.00 | 32.28 0.00 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 95.66 0.08 | 54.11 0.98 | 69.12 0.83 | 24.30 0.38 | 85.71 0.00 | 37.86 0.47 | 53.49 0.65 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 100.00 0.00 | 1.10 0.25 | 2.18 0.48 | 14.80 0.03 | 100.00 0.00 | 25.78 0.05 | 13.98 0.26 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 92.98 0.04 | 40.61 0.25 | 56.53 0.24 | 19.20 0.06 | 82.14 0.00 | 31.12 0.08 | 43.83 0.16 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 88.97 0.24 | 90.06 0.25 | 89.51 0.24 | 37.69 1.54 | 35.00 1.43 | 36.30 1.48 | 62.90 0.86 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 100.00 0.00 | 2.45 0.00 | 4.79 0.00 | 14.97 0.00 | 100.00 0.00 | 26.05 0.00 | 15.42 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 96.88 0.00 | 57.06 0.00 | 71.81 0.00 | 26.32 0.00 | 89.29 0.00 | 40.65 0.00 | 56.23 0.00 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 90.94 0.06 | 30.80 0.25 | 46.01 0.28 | 16.94 0.05 | 82.14 0.00 | 28.08 0.07 | 37.05 0.18 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 84.93 0.00 | 38.04 0.00 | 52.54 0.00 | 14.41 0.00 | 60.71 0.00 | 23.29 0.00 | 37.92 0.00 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 85.25 0.03 | 99.26 0.25 | 91.72 0.12 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 45.86 0.06 | |||||||||
| A-all | Gemma2 2B IT | 5-shot MMLU | 62.86 0.71 | 6.67 1.84 | 11.99 2.92 | 46.86 0.22 | 95.47 1.07 | 62.86 0.04 | 37.42 1.44 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 66.67 0.00 | 11.49 0.00 | 19.61 0.00 | 47.62 0.00 | 93.33 0.00 | 63.06 0.00 | 41.34 0.00 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 57.52 0.19 | 63.45 5.06 | 60.24 2.04 | 51.94 1.03 | 45.60 4.80 | 48.36 2.58 | 54.30 0.27 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 63.85 0.77 | 38.16 0.46 | 47.77 0.58 | 51.09 0.36 | 74.93 0.53 | 60.76 0.43 | 54.26 0.50 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 65.42 1.30 | 45.29 5.52 | 53.28 3.06 | 53.22 0.77 | 72.00 5.33 | 61.08 1.61 | 57.18 0.73 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 62.67 0.06 | 64.83 1.38 | 63.72 0.65 | 57.51 0.47 | 55.20 1.07 | 56.32 0.32 | 60.02 0.17 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 80.51 5.64 | 11.26 0.46 | 19.76 0.88 | 48.46 0.41 | 96.80 1.07 | 64.59 0.60 | 42.18 0.74 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 67.91 0.93 | 33.56 0.46 | 44.92 0.62 | 51.43 0.34 | 81.60 0.53 | 63.09 0.41 | 54.01 0.51 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 56.68 1.42 | 80.00 8.28 | 66.10 1.52 | 54.26 2.13 | 28.53 12.27 | 35.58 14.09 | 50.84 6.29 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 64.91 1.21 | 22.07 0.46 | 32.93 0.34 | 48.79 0.16 | 86.13 1.07 | 62.29 0.41 | 47.61 0.04 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 62.12 4.24 | 22.07 4.14 | 32.16 4.54 | 48.08 0.10 | 83.73 4.80 | 61.04 1.30 | 46.60 1.62 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 59.18 0.00 | 33.33 0.00 | 42.65 0.00 | 48.67 0.00 | 73.33 0.00 | 58.51 0.00 | 50.58 0.00 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 53.67 0.50 | 97.93 1.84 | 69.32 0.06 | 11.67 23.33 | 1.87 3.73 | 3.22 6.44 | 36.27 3.19 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 53.42 0.00 | 98.85 0.00 | 69.35 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.68 0.00 |